OpenStudy (anonymous):

Factor completely: 2x2 + 5x + 2 Prime (2x + 2)(x + 1) (2x − 2)(x − 1) (2x + 1)(x + 2)

OpenStudy (math&ing001):

When you have choices, it's easy to just distribute them and see if any matches 2x2 + 5x + 2

OpenStudy (anonymous):

Im not good with this lesson can you show mw how to do it ?

OpenStudy (math&ing001):

Alright, I'll do the first one: (2x + 2)(x + 1) = 2x2 + 4x + 2 So it's not option 2

OpenStudy (anonymous):

I don't get how to do it though

OpenStudy (math&ing001):

You can't answer this kind of questions if you don't know how to distribute. I'll try my best to explain. It works like this: a*(b + c) = a*b + a*c (a + b)(c + d) = ac + ad + bc + bd Because + is distributive !